SO,, 13.4 g 52T At a certain temperature, a 24.5-L container holds four gases in equilibrium. Their masses are 3.5 g SO2,4 NyAt a certain temperature, a 24.5-L container holds four gases in equilibrium. Their masses are 3.5 g SO3 , 4.6 g SO2 , 13.4 g N2 , and 0.98 g N2O .
Answer #1

Reaction between SO2 and N2O to form SO3 and N2 is as shown

SO2(g) + N20(9) + S03(9) + N2(9

This is the balanced chemical reaction.

Molar mass of SO2=Molar mass of S+2xMolar mass of O=32 g/mol+2x16 g/mol=32 g/mol+32 g/mol=64 g/mol

Molar mass of N2O=2xMolar mass of N+Molar mass of O=2x14 g/mol+16 g/mol=28 g/mol+16 g/mol=44 g/mol

Molar mass of SO3=Molar mass of S+3xMolar mass of O=32 g/mol+3x16 g/mol=32 g/mol+48 g/mol=80 g/mol

Molar mass of N2=2xMolar mass of N=2x14 g/mol=28 g/mol

Number of moles of SO2=Given mass/Molar mass=4.6 g/64 g/mol=0.07 mol

Number of moles of N2O=Given mass/Molar mass=0.98 g/44 g/mol=0.02 mol

Number of moles of SO3=Given mass/Molar mass=3.5 g/80 g/mol=0.04 mol

Number of moles of N2=Given mass/molar mass=13.4 g/28 g/mol=0.48 mol

We know that concentration=number of moles/volume of solution (L)

Also equilibrium constant Kc is the ratio of product of concentrations of products at equilibrium to product of concentrations of reactants at equilibrium raised to power their stoichiometric coefficients.

So Kc=[SO3][N2]/[SO2][N2O]

=(0.04 mol/24.5 L)(0.48 mol/24.5 L)/{(0.07 mol/24.5 L)(0.02 mol/24.5 L)}

=13.71

So the equilibrium constant Kc=13.71